OpenStudy (anonymous):

i need help on another problem a fraction into a mixed number... it's .... 14 5/6 - 7 11/12=

OpenStudy (anonymous):

come to http://www.dabbleboard.com/draw?b=Guest618759&i=0&c=47eb51b00443cd8511819e96601d62f3ed7343b7 I am helping someone else there right now too but I can help you

OpenStudy (anonymous):

14 5/6=178/12 7 11/12=95/12 subtract 95 from 178 you have 83 left so it is 83/12 or broken down is 6 11/12

OpenStudy (anonymous):

common denominator for the two numbers is twelve do you know how I got that no offense but this is simple math